A total of $7100 was invested as follows: a certain amount at 7%, twice that amount at 10%, and the remainder at 12%. If the annual interest from the three investments was $699, how much was invested at each rate?
Answer by mananth(16946)

A total of $7100 was invested .
a certain amount at 7%,-----------x
twice that amount at 10%,----------2x
the remainder at 12%. ---------------(7100-3x)
annual interest from the three investments = $699
7%x+10%*2x+12%*(7100-3x)=699
multiply by 100
7x+20x+12(7100-3x)=699
7x+20x+85200-36x=699
27x-36x=699-85200
-9x=-84501
/-9
x=9389
amount at 7%,-----------$9389
twice that amount at 10%,----------2x---------------$18778
the remainder at 12%. --------------- Balance